# PlainQAFact: Automatic Factuality Evaluation Metric for Biomedical Plain Language Summaries Generation

### v1

@@ Appendix G Pilot Study on FactPICO
In PLS generation tasks, we employ the recently introduced FactPICO dataset Joseph et al. (2024). This dataset comprises human-labeled plain language summaries of Randomized Controlled Trials (RCTs) that address several critical elements: Populations, Interventions, Comparators, Outcomes (PICO), and additional information. All summaries are generated by various LLMs based on scientific literature and include extensive explanations not present in the original abstracts. We hypothesize that existing factuality evaluation metrics for text summarization may struggle to accurately assess the factuality of this added information. To validate this assumption, we conduct pilot studies using four factuality evaluation metrics from FactPICO alongside one NLI-based metric: DAE Goyal and Durrett (2021), AlignScore Zha et al. (2023), SummaC Laban et al. (2022), QAFactEval Fabbri et al. (2022a), and QuestEval Scialom et al. (2021b).
@@ G.2 Experiments and Results
We conduct experiments using the processed FactPICO dataset to address both research questions. It is important to note that there are some outliers in the PLSs which contain extraneous content and exceed the input length limitations of the DAE and QAFactEval metrics. Therefore, we are unable to include these two metrics in this pilot study.
Figure 6: Score change percentage from baselines over three metrics on the FactPICO dataset in removing factual added information (88 pairs). We expect each model get higher performance when more external information is removed. The evaluation dataset contains 88 valid summary-abstract pairs.
RQ1: Do existing metrics perform well when external information is added to PLS compared with no added information?
In this study, we evaluate the ability of existing metrics to detect added information in PLSs. We focus on PLSs where all added information is annotated as factual, resulting in a deduplicated dataset of 88 PLS-abstract pairs. To assess metric sensitivity to the added information, we iteratively remove sentences from each PLS that contain added spans. For example, if a span such as “of a medicine called haloperidol” is labeled as factual (“yes”), we remove the entire sentence in the original PLS containing that span through exact matching, continuing this process until no added information remains. This procedure allows us to determine how effectively existing metrics handle external information that is not present in the original abstracts. As shown in Figure 6, we report performance changes relative to baseline scores. For example, AlignScore increases by 3.33 (on a scale of 100) when six units of added information are removed, leading to a change of approximately 4.9%. These results indicate that added information has impacts on model-based factuality evaluation metrics such as AlignScore and SummaC, with both scores improving as more added information is removed. In contrast, QuestEval’s score decreases with the removal of added information.
Figure 7: Score change percentage from baselines over three metrics on the FactPICO dataset in removing non-factual added information (60 pairs). We expect each metric’s score increases when more non-factual information is removed.
@@
In this research question, our goal is to evaluate the sensitivity of factuality evaluation metrics in detecting non-factual information within PLSs. The FactPICO dataset labels added information as either “yes” (factual) or “no” (non-factual). For this scenario, we sample only those PLSs that contain both “yes” and “no” labels for added spans. As with RQ1, we iteratively remove sentences containing non-factual added spans from each PLS until no non-factual sentences remain. Overall, as illustrated in Figure 7, most evaluation metrics show improved performance as more non-factual information is removed. Nevertheless, our findings suggest that existing factuality evaluation metrics have limited capacity to accurately distinguish between factual and non-factual added information in text summarization tasks.

### v2

@@ Appendix G Pilot Study on FactPICO
To investigate the performance of existing automatic factual consistency evaluation metrics on plain language generation tasks, we employ the recently introduced FactPICO dataset [9]. This dataset comprises human-labeled plain language summaries of Randomized Controlled Trials (RCTs) that address several critical elements: Populations, Interventions, Comparators, Outcomes (PICO), and additional information. All summaries are generated by various LLMs based on medical literature and include added information (i.e., extensive explanations) not present in the original abstracts. We hypothesize that existing factuality evaluation metrics for text summarization may struggle to accurately assess the factuality of this added information. To validate this assumption, we conduct pilot studies using four factual consistency evaluation metrics from FactPICO alongside one NLI-based metric: DAE [48], AlignScore [19], SummaC [16], QAFactEval [39], and QuestEval [50].
@@ G.2 Experiments and Results
We conduct experiments using the processed FactPICO dataset to conduct pilot studies that assess the performance of existing factual consistency evaluation metrics in detecting added information in plain language summarization tasks. It is important to note that some outliers in the plain language summaries of the FactPICO dataset contain extraneous content, which causes the input lengths to exceed the limitations of both the DAE and QAFactEval metrics. Furthermore, the factuality of the LLM-generated plain language summaries is not guaranteed.
Figure 5: Score change percentage from baselines over five metrics on the FactPICO dataset in removing factual added information. We expect each metric stays unchanged even when more added factual information is removed. The evaluation dataset contains 88 valid summary-abstract pairs.
Figure 6: Score change percentage from baselines over five metrics on the FactPICO dataset in removing non-factual added information (60 pairs). We expect the change percentage from baseline increases when more added non-factual information is removed.
RQ1: Do existing metrics perform well when external information is added to plain language summaries compared with no added information?
In this study, we evaluate the ability of existing factual consistency evaluation metrics to detect added information in plain language summaries. We focus on summaries where all added information is annotated as factual, resulting in a dataset of 88 summary-abstract pairs. To assess metric sensitivity to the added information, we iteratively remove sentences from each plain language summary that contain added spans. For example, if a span such as “of a medicine called haloperidol” is labeled as factual (“yes”), we remove the entire sentence in the original plain language summaries containing that span through exact matching, continuing this process until no added information remains. This procedure enables us to determine how effectively current metrics handle added information that is absent from the original abstracts. As shown in Figure 5, we report performance changes relative to baseline scores. For example, the AlignScore evaluation increases by 3.3% (on a 100-point scale) when six spans of added information are removed, resulting in an overall change of approximately 4.9% compared to the setting in which no added information is removed. These findings indicate that added information affects model-based factuality evaluation metrics such as QAFactEval (6.0%), AlignScore (4.9%), and SummaC (1.9%), with scores improving as more added information is removed. In contrast, both QuestEval and DAE scores decline with the removal of added information, and notably, DAE exhibits the most significant performance drop. Overall, these findings suggest that all the five metrics have difficulty accurately assessing added factual information, as evidenced by both increases and decreases in their performance.
@@
In this research question, our goal is to evaluate the sensitivity of factuality evaluation metrics in detecting non-factual information within plain language summaries. The FactPICO dataset labels added information as either “yes” (factual) or “no” (non-factual). In this scenario, we sample only those plain language summaries that contain both “yes” and “no” labels for added spans. As with RQ1, we iteratively remove sentences containing non-factual added spans from each plain language summary until no non-factual sentences remain. Overall, as illustrated in Figure 6, only AlignScore, SummaC, and QAFactEval show improved performance as more non-factual information is removed, indicating that these metrics are sensitive in added non-factual information. Nevertheless, based on the results of RQ 1 and 2, our findings suggest that existing factual consistency evaluation metrics have limited capacity to accurately distinguish between factual and non-factual added information in plain language summarization tasks.